a paallelogram with sides 12 and 16 and a angle of the measure 60. i need to no the area of it ?
Answer by jim_thompson5910(35256)

Area = ab*sin(theta)

A = 12*16*sin(60)

A = 192*sin(60)

A = 192*sqrt(3)/2

A = 96*sqrt(3)

So the exact area is 96*sqrt(3) square units

Making the approximate area to be 166.2768775 square units